# Simultaneous quantum identity authentication scheme utilizing entanglement swapping with secret key preservation

### v1

@@ I Introduction
The first QIA scheme was introduced by Crépeau et al. [14] in 1995, utilizing oblivious transfer (OT) as a cryptographic primitive. However, subsequent work by Lo and Chau [28] demonstrated that quantum OT cannot achieve unconditional security in a two-party scenario. Consequently, Crépeau et al.’s QIA scheme lacked unconditional security. It is noteworthy that OT is not the sole cryptographic task applicable to QIA scheme design. Following Crépeau et al.’s work, a series of QIA protocols emerged [15, 16, 17, 18, 19, 20, 21, 29, 30, 22]. The schemes for QIA are built upon protocols designed for various cryptographic tasks, serving as the foundation for their development. For a comprehensive understanding of QIA and the adaptation of diverse cryptographic tasks to formulate QIA schemes, refer to the review provided in Ref. [22]. Some of these tasks were adapted from protocols for secure direct quantum communication, which involve transmitting messages using quantum resources without key generation [31, 32, 33]. Before delving into a summary of existing QIA schemes derived from modified secure direct communication schemes, it is pertinent to distinguish between two types of secure quantum direct communication: quantum secure direct communication (QSDC) and deterministic secure quantum communication (DSQC). Numerous QSDC and DSQC schemes exist [31, 34, 35, 36, 37, 38, 39, 40], with many early ones being entangled state-based [31, 38, 41]. For instance, Zhang et al. [17] adapted the entangled state-based ping-pong protocol [41] for QSDC to devise a QIA scheme. Moreover, Yuan et al. [42] proposed a QIA protocol based on the LM05 protocol [32], which is the single photon analog of the ping-pong protocol. Other authors have also modified DSQC and QSDC protocols to create QIA schemes [15, 43, 44]. Additionally, novel schemes have emerged for controlled DSQC (CDSQC) [35, 45], where a semi-honest Alice securely communicates with Bob using quantum resources only under the supervision of a controller, Charlie. The potential of adapting CDSQC schemes for QIA remains relatively unexplored. Taking into account the significance of identity authentication, we aim to propose a new QIA protocol with simultaneous authentication of legitimate parties, inspired by CDSQC principles. In the following sections, we will introduce a Bell state-based protocol for QIA, drawing from CDSQC concepts, and demonstrate its resilience against common attacks.
The remainder of this letter is structured as follows: Section II introduces the proposed QIA protocol that uses Bell states. Section III critically analyzes the protocol’s security, demonstrating its robustness against multiple types of attacks. Section IV compares our protocol with the existing ones from the same family. Finally, Section V concludes the letter.
@@ III Evaluation of the security aspects in the proposed protocol
In this section, we aim to evaluate the security of the proposed protocol against various well-known attacks that can be attempted by an eavesdropper, commonly referred to as Eve. It is important to mention that insider attacks (by Alice and Bob) will not be discussed here, as they are not pertinent within the scope of QIA. Initially, we will focus on the impersonation attack. In this scenario, we consider the possibility of an untrusted participant, Charlie, attempting to gain unauthorized access to confidential information while adhering to all steps of the protocol. Moreover, an impersonated fraudulent attack entails an external party, such as Eve, aiming to mimic the legitimate user, either Alice or Bob, and effectively maneuver through the authentication procedure.
@@ III.2 Security assessment of the protocol against intercept and resend attack
In this quantum protocol, Alice and Bob generate Bell states based on their pre-shared secret key. Instead of transmitting transmitting their entire quantum states, Alice and Bob selectively send Particle 2 and Particle 4 through the quantum channel. This approach limits Eve’s potential attacks to these specific particles at any given moment. For the sake of simplicity, we assume that Eve will attempt to intercept Particle 2 and Particle 4 as they are transmitted from Alice to Charlie and from Bob to Alice, respectively. Eve’s ability to extract information from the quantum channel is constrained by the Holevo bound or Holevo quantity [49]
@@
where $\rho^{24}$ and $\rho_{i}^{24}$ represent the reduced matrices of $\rho$ and $\rho_{i}$, respectively, obtained after performing a partial trace over particles 1 and 3. Through straightforward calculations, we determine that $\rho^{24}={\rm Tr}_{13}\left(\rho\right)=\frac{1}{4}\left(|11\rangle\langle 11|+|10\rangle\langle 10|+|01\rangle\langle 01|+|00\rangle\langle 00|\right)=\frac{1}{4}\mathds{1}_{4}$, where $\mathds{1}_{4}$ denotes the $4\times 4$ identity matrix. Similarly, the von Neumann entropy of each component of the mixed state $\rho$ after the partial trace, denoted as $\rho_{i}^{24}={\rm Tr}_{13}\left(\rho_{i}\right)=\frac{1}{4}\mathds{1}_{4}$. By substituting $\rho^{24}$ and $\rho_{i}^{24}$ into Eq. (8), we deduce that $\chi\left(\rho^{24}\right)=0$. Consequently, we infer that Eve cannot obtain any key information through direct intercept attacks on the transmission particles. The intercept-resend (IR) attack strategy employed by Eve is depicted in Fig. 3.
@@ III.3 Security assessment of the protocol against impersonated fraudulent attack
IV Comparison with the existing protocols
@@ V Conclusion
The present work, along with references [22, 23], suggests the potential for developing a wide range of new QIA schemes. A comprehensive exploration in this direction can be highly valuable for identifying the most efficient scheme for QIA that can be implemented using current technology. Our proposed protocol for QIA is an entangled-state-based scheme for QIA that utilizes Bell states. Inspired by the principles of secure direct quantum communication, it enables Alice to securely transmit information to Bob when controller Charlie authorizes it, without the need for pre-generated keys. Although the protocol relies solely on Bell states, its realization would necessitate quantum memory, which is not yet commercially available. While this represents a limitation, it is not unique to our protocol, but this limitation is common to many existing protocols for QIA, QSDC, DSQC and quantum dialogue. Given the growing interest in research and technology development related to quantum memory, along with recent proposals for constructing quantum memory, it is anticipated that quantum memory will become available in the near future [54]. In the interim period, a delay mechanism can serve as a substitute for quantum memory in implementing our proposed QIA protocol, thereby enhancing the resilience of protocols for secure quantum communication.

### v2

@@ I Introduction
The first QIA scheme was introduced by Crépeau et al. [14] in 1995, utilizing oblivious transfer (OT) as a cryptographic primitive. However, subsequent work by Lo and Chau [28] demonstrated that quantum OT cannot achieve unconditional security in a two-party scenario. Consequently, Crépeau et al.’s QIA scheme lacked unconditional security. It is noteworthy that OT is not the sole cryptographic task applicable to QIA scheme design. Following Crépeau et al.’s work, a series of QIA protocols emerged [15, 16, 17, 18, 19, 20, 21, 29, 30, 22]. The schemes for QIA are built upon protocols designed for various cryptographic tasks, serving as the foundation for their development. For a comprehensive understanding of QIA and the adaptation of diverse cryptographic tasks to formulate QIA schemes, refer to the review provided in Ref. [22]. Some of these tasks were adapted from protocols for secure direct quantum communication, which involve transmitting messages using quantum resources without key generation [31, 32, 33]. Before delving into a summary of existing QIA schemes derived from modified secure direct communication schemes, it is pertinent to distinguish between two types of secure quantum direct communication: quantum secure direct communication (QSDC) and deterministic secure quantum communication (DSQC). Numerous QSDC and DSQC schemes exist [31, 34, 35, 36, 37, 38, 39, 40], with many early ones being entangled state-based [31, 38, 41]. For instance, Zhang et al. [17] adapted the entangled state-based ping-pong protocol [41] for QSDC to devise a QIA scheme. Moreover, Yuan et al. [42] proposed a QIA protocol based on the LM05 protocol [32], which is the single photon analog of the ping-pong protocol. Other authors have also modified DSQC and QSDC protocols to create QIA schemes [15, 43, 44]. Additionally, novel schemes have emerged for controlled DSQC (CDSQC) [35, 45], where a semi-honest Alice securely communicates with Bob using quantum resources only under the supervision of a controller, Charlie. The potential of adapting CDSQC schemes for QIA remains relatively unexplored. Taking into account the significance of identity authentication, we aim to propose a new QIA protocol with simultaneous authentication of legitimate parties, inspired by CDSQC principles. In the following sections, we will introduce a Bell state-based protocol for QIA, drawing from CDSQC concepts, and demonstrate its resilience against common attacks. Additionally, we will analyze the impact of collective noise on the proposed scheme.
The remainder of this letter is structured as follows: Section II introduces the proposed QIA protocol that uses Bell states. Section III critically analyzes the protocol’s security, demonstrating its robustness against multiple types of attacks. Section IV focuses on collective noise analysis for the new QIA scheme. Section V compares our protocol with the existing ones from the same family. Finally, Section VI concludes the letter.
@@ III Evaluation of the security aspects in the proposed protocol
In this section, we aim to evaluate the security of the proposed protocol against various well-known attacks that can be attempted by an eavesdropper, commonly referred to as Eve. It is important to mention that insider attacks (by Alice and Bob) will not be discussed here, as they are not pertinent within the scope of QIA. Initially, we will focus on the impersonation attack. In this scenario, we consider the possibility of an untrusted participant, Charlie, attempting to gain unauthorized access to confidential information while adhering to all steps of the protocol. Moreover, an impersonated fraudulent attack entails an external party, such as Eve, aiming to mimic the legitimate user, either Alice or Bob, and effectively maneuver through the authentication procedure. In particular, among the attacks discussed in this section, the intercept-resend attack and impersonation-based fraudulent attack specifically target the quantum channel.
@@ III.2 Security assessment of the protocol against intercept and resend attack
In this quantum protocol, Alice and Bob generate Bell states based on their pre-shared secret key. Instead of transmitting their entire quantum states, Alice and Bob selectively send Particle 2 and Particle 4 through the quantum channel. This approach limits Eve’s potential attacks to these specific particles at any given moment. For the sake of simplicity, we assume that Eve will attempt to intercept Particle 2 and Particle 4 as they are transmitted from Alice to Charlie and from Bob to Alice, respectively. Eve’s ability to extract information from the quantum channel is constrained by the Holevo bound or Holevo quantity [49]
@@
where $\rho^{24}$ and $\rho_{i}^{24}$ represent the reduced matrices of $\rho$ and $\rho_{i}$, respectively, obtained after performing a partial trace over particles 1 and 3. Through straightforward calculations, we determine that $\rho^{24}={\rm Tr}_{13}\left(\rho\right)=\frac{1}{4}\left(|11\rangle\langle 11|+|10\rangle\langle 10|+|01\rangle\langle 01|+|00\rangle\langle 00|\right)=\frac{1}{4}\mathds{1}_{4}$, where $\mathds{1}_{4}$ denotes the $4\times 4$ identity matrix. Similarly, the von Neumann entropy of each component of the mixed state $\rho$ after the partial trace, denoted as $\rho_{i}^{24}={\rm Tr}_{13}\left(\rho_{i}\right)=\frac{1}{4}\mathds{1}_{4}$. By substituting $\rho^{24}$ and $\rho_{i}^{24}$ into Eq. (8), we deduce that $\chi\left(\rho^{24}\right)=0$. Consequently, we infer that Eve cannot obtain any key information through direct intercept attacks on the transmission particles. The intercept-resend (IR) attack strategy employed by Eve is depicted in Fig. 3. The maximally mixed state is considered with equal probabilities of Bell states in Eq. (7) to minimize or eliminate Eve’s information gain from the channel in IR attack. If these probabilities are not equal, the security against the IR attack will depend on the specific values of the Bell state probabilities in the mixed state. As long as the mutual information between Alice and Bob remains greater than the Holevo information $\left(I(A|B)>\chi\left(\rho\right)\right)$, the scheme will remain secure against the IR attack [50].
@@ III.3 Security assessment of the protocol against impersonated fraudulent attack
IV Collective noise analysis of proposed protocol
The QIA protocol mentioned earlier was discussed assuming an ideal quantum channel. In practice, when particles are transmitted through a quantum channel, they are exposed to noise, which can alter their state(s). Additionally, an attacker might exploit the presence of noise to mask his attack(s), making it difficult to distinguish between errors caused by noise and those introduced by an attacker. Addressing the impact of collective noise on quantum communication is a well-known challenge [51, 52, 53]. Walton et al. [54] introduced the concept of a decoherence-free subspace (DFS), which can mitigate the effects of collective noise since the states of the particles remain unaffected within these channels. Further research has also been conducted on quantum communication under collective noise [55, 56]. Collective noise typically includes both collective-dephasing and collective-rotation noise [57]. In what follows, we will analyze the impact of these types of noise on the scheme proposed here.
IV.1 Collective-dephasing noise
A collective-dephasing noise can be expressed as [57, 56]
$\begin{array}[]{lcl}U_{dp}|0\rangle=|0\rangle&,&U_{dp}|1\rangle=e^{i\phi}|1\rangle\end{array}.$
Collective-dephasing noise is characterized by a parameter $\phi$, which varies with time. Typically, a logical qubit encoded in the product states of two physical qubits as
$\begin{array}[]{lclc}|0\rangle_{L}=|01\rangle&,&|1\rangle_{L}=|10\rangle&,\end{array}$
is resistant to this type of noise because both logical qubits accumulate the same phase factor $e^{i\phi}$.
In this context, the subscript $L$ denotes the logical qubit, while 0 and 1 correspond to horizontal and vertical polarization states, respectively; these being the eigenstates of the Pauli operator $\sigma_{z}$ ($Z$ basis). The states $|+_{L}\rangle$ and $|-_{L}\rangle$ are defined as follows:
$\begin{array}[]{ccccc}|+_{L}\rangle&=&\frac{1}{\sqrt{2}}\left(|0_{L}\rangle+|1_{L}\rangle\right)&=&\frac{1}{\sqrt{2}}\left(|01\rangle+|10\rangle\right)\\ |-_{L}\rangle&=&\frac{1}{\sqrt{2}}\left(|0_{L}\rangle-|1_{L}\rangle\right)&=&\frac{1}{\sqrt{2}}\left(|01\rangle+|10\rangle\right)\end{array}.$
From the equations presented in Section II.1, specifically from (3) and (6), it is evident that particles 2 and 4 in the composite states are responsible for transmitting quantum information through the quantum channel. Taking Eq. (3) as an example:
$\begin{array}[]{lcl}\left|\Psi^{11}\right\rangle&=&\sigma_{z1}\otimes\sigma_{z3}\left(\left|\phi^{+}\right\rangle_{12}\left|\psi^{-}\right\rangle_{34}\right)\\ &=&\frac{1}{2}\sigma_{z1}\otimes\sigma_{z3}\left(|0001\rangle-|0010\rangle+|1101\rangle-|1110\rangle\right)_{1234}\\ &=&\frac{1}{2}\left(|0001\rangle+|0010\rangle-|1101\rangle-|1110\rangle\right)_{1234}\end{array},$
after considering collective-dephasing noise on particles $2$ and $4$, we obtain
$\begin{array}[]{lcl}\left|\Psi^{11}\right\rangle_{dp}&=&\frac{1}{2}\left(e^{i\phi}|0001\rangle+|0010\rangle-e^{2i\phi}|1101\rangle-e^{i\phi}|1110\rangle\right)_{1234}\\ &=&\frac{1}{2}\left(e^{i\phi}|01\rangle|00\rangle+|00\rangle|01\rangle-e^{2i\phi}|11\rangle|10\rangle-e^{i\phi}|10\rangle|11\rangle\right)_{1423}\\ &=&\frac{1}{4}\left[2e^{i\phi}\left(|\psi^{+}\rangle|\phi^{-}\rangle+|\psi^{-}\rangle|\phi^{+}\rangle\right)+\left(1-e^{2i\phi}\right)\left(|\phi^{+}\rangle|\psi^{+}\rangle+|\phi^{-}\rangle|\psi^{-}\rangle\right)_{1423}\right..\\ &+&\left.\left(1+e^{2i\phi}\right)\left(|\phi^{+}\rangle|\psi^{-}\rangle+|\phi^{-}\rangle|\psi^{+}\rangle\right)\right]_{1423}\end{array}$
[…]
Figure 4: The collective error probability with respect to the noise parameter can be categorized as: (a) Collective dephasing error probability with noise parameter $\phi$, and (b) Collective rotation error probability with noise parameter $\theta$.
IV.2 Collective-rotation noise
The collective-rotation noise can be represented as:
$\begin{array}[]{lcl}U_{r}|0\rangle=\cos\theta|0\rangle+\sin\theta|1\rangle&,&U_{r}|1\rangle=-\sin\theta|0\rangle+\cos\theta|1\rangle\end{array},$
The collective rotation noise is represented by the parameter $\theta$, which varies over time in the quantum channel. For example, applying collective rotation noise to particles 2 and 4 from Eq. (3) results in the following expression after a tedious calculations,
$\begin{array}[]{lcl}U_{r2}\otimes U_{r4}\left|\Psi^{11}\right\rangle&=&\frac{1}{2}\left[-\sin\theta\cos\theta|0000\rangle-\sin^{2}\theta|0010\rangle+\cos^{2}\theta|0100\rangle+\sin\theta\cos\theta|0110\rangle\right.\\ &+&\sin\theta\cos\theta|1001\rangle-\cos^{2}\theta|1011\rangle+\sin^{2}\theta|1101\rangle-\sin\theta\cos\theta|1111\rangle\\ &-&\left(\cos^{2}\theta|0001\rangle+\sin\theta\cos\theta|0011\rangle+\sin\theta\cos\theta|0101\rangle+\sin^{2}\theta|0111\rangle\right.\\ &-&\left.\left.\sin^{2}\theta|1000\rangle+\sin\theta\cos\theta|1010\rangle+\sin\theta\cos\theta|1100\rangle-\cos^{2}\theta|1110\rangle\right)\right]_{1423}.\\ &=&\frac{1}{2}\left[2\sin\theta\cos\theta\left(-|\phi^{+}\rangle|\phi^{+}\rangle-|\psi^{-}\rangle|\psi^{-}\rangle\right)\right.\\ &+&\sin^{2}\theta\left(|\phi^{+}\rangle|\psi^{-}\rangle-|\phi^{-}\rangle|\psi^{+}\rangle+|\psi^{+}\rangle|\phi^{-}\rangle-|\psi^{-}\rangle|\phi^{+}\rangle\right)\\ &+&\left.\cos^{2}\theta\left(|\psi^{+}\rangle|\phi^{-}\rangle+|\psi^{-}\rangle|\phi^{+}\rangle-|\phi^{+}\rangle|\psi^{-}\rangle-|\phi^{-}\rangle|\psi^{+}\rangle\right)\right]_{1423}\end{array}$
In the absence of noise or errors, the Bell pairs shared by Alice and Bob are $\left\{|\psi^{-}\rangle|\phi^{+}\rangle,|\psi^{+}\rangle|\phi^{-}\rangle,|\phi^{-}\rangle|\psi^{+}\rangle,|\phi^{+}\rangle|\psi^{-}\rangle\right\}$. From this above result, the probability of error due to collective rotation noise can be derived as $2\sin^{2}\theta\cos^{2}\theta$, which depends on the noise parameter $\theta$. Similarly, applying collective rotation noise to particles 2 and 4 from Eq. (4) yields:
$\begin{array}[]{lcl}U_{r2}\otimes U_{r4}\left|\Psi^{00}\right\rangle&=&\frac{1}{2}\left[\cos^{2}\theta|0000\rangle+\sin\theta\cos\theta|0010\rangle+\sin\theta\cos\theta|0100\rangle+\sin^{2}\theta|0110\rangle\right.\\ &+&\sin^{2}\theta|1001\rangle-\sin\theta\cos\theta|1011\rangle-\sin\theta\cos\theta|1101\rangle+\cos^{2}\theta|1111\rangle\\ &-&\left(-\sin\theta\cos\theta|0001\rangle-\sin^{2}\theta|0011\rangle+\cos^{2}\theta|0101\rangle+\sin\theta\cos\theta|0111\rangle\right.\\ &-&\left.\left.-\sin\theta\cos\theta|1000\rangle+\cos^{2}\theta|1010\rangle-\sin^{2}\theta|1100\rangle+-\sin\theta\cos\theta|1110\rangle\right)\right]_{1423}.\\ &=&\frac{1}{2}\left[2\sin\theta\cos\theta\left(|\phi^{-}\rangle|\psi^{+}\rangle+|\psi^{+}\rangle|\phi^{-}\rangle\right)\right.\\ &+&\sin^{2}\theta\left(|\psi^{+}\rangle|\psi^{+}\rangle-|\psi^{-}\rangle|\psi^{-}\rangle+|\phi^{+}\rangle|\phi^{+}\rangle-|\phi^{-}\rangle|\phi^{-}\rangle\right)\\ &+&\left.\cos^{2}\theta\left(|\phi^{+}\rangle|\phi^{+}+|\phi^{-}\rangle|\phi^{-}\rangle+|\psi^{+}\rangle|\psi^{+}\rangle+|\psi^{-}\rangle|\psi^{-}\rangle\right)\right]_{1423}\end{array}$
[…]
V Comparison with the existing protocols
@@ VI Conclusion
The present work, along with references [22, 23], suggests the potential for developing a wide range of new QIA schemes. A comprehensive exploration in this direction can be highly valuable for identifying the most efficient scheme for QIA that can be implemented using current technology. Our proposed protocol for QIA is an entangled-state-based scheme for QIA that utilizes Bell states. Inspired by the principles of secure direct quantum communication, it enables Alice to securely transmit information to Bob when controller Charlie authorizes it, without the need for pre-generated keys. Although the protocol relies solely on Bell states, its realization would necessitate quantum memory, which is not yet commercially available. While this represents a limitation, it is not unique to our protocol, but this limitation is common to many existing protocols for QIA, QSDC, DSQC and quantum dialogue. Given the growing interest in research and technology development related to quantum memory, along with recent proposals for constructing quantum memory, it is anticipated that quantum memory will become available in the near future [65]. In the interim period, a delay mechanism can serve as a substitute for quantum memory in implementing our proposed QIA protocol, thereby enhancing the resilience of protocols for secure quantum communication. Additionally, our work could be extended in the future with modifications that provide resistance to collective noise [58, 59, 60].